### v1

# Character Complexity: A Novel Measure for Quantum Circuit Analysis

Abstract
In the rapidly evolving field of quantum computing, quantifying circuit complexity remains a critical challenge. This paper introduces character complexity, a novel measure that bridges group-theoretic concepts with practical quantum computing concerns. By leveraging tools from representation theory, I prove several key properties of character complexity and establish a surprising connection to the classical simulability of quantum circuits. This new measure offers a fresh perspective on the complexity landscape of quantum algorithms, potentially reshaping our understanding of quantum-classical computational boundaries. I present innovative visualization methods for character complexity, providing intuitive insights into the structure of quantum circuits. The empirical results reveal intriguing scaling behaviors with respect to qubit and gate counts, opening new avenues for quantum algorithm design and optimization. This work not only contributes to the theoretical foundations of quantum complexity but also offers practical tools for the quantum computing community. As quantum hardware continues to advance, character complexity could play a crucial role in developing more efficient quantum algorithms and in exploring the fundamental limits of quantum computation.
1 Introduction
Quantum computing has witnessed remarkable progress in recent years, with the development of increasingly sophisticated quantum algorithms and the rapid advancement of quantum hardware. However, as the complexity of these algorithms grows, we face a critical challenge in efficiently analyzing and optimizing them. Traditional complexity measures, such as circuit depth and gate count, while valuable, often fail to capture the intricate mathematical structure that underlies the power of quantum computation. This limitation hinders our ability to fully harness the potential of quantum algorithms and to push the boundaries of quantum-classical computational advantage.
In this paper, I introduce a groundbreaking approach to quantifying the complexity of quantum circuits: Character Complexity. This novel measure, rooted in the rich mathematical framework of group representation theory, offers an unprecedented lens through which to analyze and understand the behavior of quantum algorithms. By leveraging the power of character theory, a fundamental tool in representation theory, I uncover hidden symmetries and structures within quantum circuits that are invisible to conventional complexity measures.
The character complexity measure offers several significant advantages over existing metrics:
Unparalleled Insight: Character complexity captures subtle nuances in the behavior of quantum circuits that are indistinguishable under metrics like gate count or circuit depth. This measure provides a more fine-grained understanding of the complexity landscape of quantum algorithms, enabling researchers to identify and exploit structural properties that may lead to more efficient implementations.
Quantum-Classical Boundary: One of the most profound contributions of this work is a theorem that establishes a tight connection between character complexity and the classical simulability of quantum circuits. I prove that quantum circuits with character complexity bounded by $C(U)\leq c\log(n)$, where $n$ is the number of qubits and $c$ is a constant, can be efficiently simulated on classical computers. This result provides a powerful new tool for delineating the boundary between quantum and classical computation, with far-reaching implications for our understanding of quantum advantage.
Visualization and Intuition: To facilitate a deeper understanding of character complexity, I introduce novel visualization techniques that provide intuitive insights into the complexity landscape of quantum circuits. These visualizations, based on a modified Bloch sphere representation, allow researchers to identify patterns, symmetries, and potential bottlenecks in quantum algorithms, guiding the design and optimization process.
The introduction of character complexity marks a significant leap forward in our understanding of the intricacies of quantum computation. By bridging the gap between the abstract mathematical world of group representation theory and the concrete realm of quantum circuits, this work provides a powerful new framework for analyzing, optimizing, and understanding the true potential of quantum algorithms. As we stand on the cusp of a new era in quantum computing, the insights and tools provided by character complexity will undoubtedly play a pivotal role in shaping the future of this exciting field.
Key Findings
  1. 1.
    Novel Complexity Measure: I introduce “Character Complexity” as a new measure for quantum circuit analysis, based on group-theoretic principles.
  2. 2.
    Theoretical Bounds: Character Complexity ($C(U)$) is bounded by $0\leq C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$, where $n$ is the number of qubits.
  3. 3.
    Composition Properties: Character Complexity is multiplicative under tensor products and submultiplicative under circuit composition.
  4. 4.
    Classical Simulation Threshold: Quantum circuits with $C(U)\leq c\log(n)$ can be efficiently simulated classically under certain conditions, providing a new criterion for classical-quantum boundaries.
  5. 5.
    Geometric Interpretation: I provide a novel visualization of Character Complexity using a modified Bloch sphere representation.
  6. 6.
    Scaling Behavior: Empirical results suggest Character Complexity scales as $O(n\log m)$ with $n$ qubits and $m$ gates, offering insights into quantum circuit structure.
  7. 7.
    Practical Implications: Character Complexity provides new perspectives on quantum algorithm design and optimization, potentially leading to more efficient quantum circuits.
2 Background
2.1 Quantum Circuits and Their Representation
Quantum circuits form the backbone of quantum algorithms, providing a structural representation of quantum computations. At their core, these circuits consist of a series of quantum gates acting on a register of qubits. While this description is intuitive, the mathematical formalism underlying quantum circuits is rich and multifaceted.
A quantum circuit acting on $n$ qubits can be represented as a unitary operator $U\in U(2^{n})$, where $U(2^{n})$ is the group of $2^{n}\times 2^{n}$ unitary matrices. This unitary nature ensures that the evolution of the quantum state remains reversible and probability-preserving. Mathematically, we can express this as:
|$\ket{\psi_{\text{out}}}=U\ket{\psi_{\text{in}}}$ (1)
where |$\ket{\psi_{\text{in}}}$ and |$\ket{\psi_{\text{out}}}$ are the input and output quantum states, respectively. While this matrix representation is complete, it often obscures the fine-grained structure of the circuit. In practice, quantum circuits are built from a set of elementary gates, each acting on a small number of qubits. Common gates include single-qubit rotations (e.g., Pauli gates, Hadamard gate) and two-qubit entangling gates (e.g., CNOT, CZ). The overall unitary $U$ is then a product of these elementary gates:
$U=U_{m}\cdots U_{2}U_{1}$ (2)
where each $U_{i}$ represents an elementary gate operation.
This decomposition into elementary gates forms the basis for many traditional complexity measures. For instance, the circuit depth is the length of the longest path through the circuit, while the T-count measures the number of T gates, which are often considered costly in fault-tolerant implementations [1]. However, these gate-based measures, while practical, often fail to capture the full quantum nature of the computation. They do not directly quantify properties such as superposition, entanglement, or the distribution of the circuit’s action across different subspaces of the Hilbert space.
An alternative representation, which forms the foundation for our character complexity measure, views quantum circuits through the lens of group theory. In this framework, we consider the group generated by the elementary gates used in the circuit. For instance, the Clifford group, generated by the Hadamard, Phase, and CNOT gates, plays a central role in many quantum error-correcting codes and can be efficiently simulated classically [2].
By representing quantum circuits as elements of finite groups, we can leverage powerful tools from representation theory, particularly the theory of group characters. This approach allows us to decompose quantum operations into irreducible representations, providing a more nuanced view of their action on the quantum state space.
It is this group-theoretic perspective that we will exploit in defining and analyzing character complexity, offering a new way to quantify the complexity of quantum circuits that goes beyond traditional gate-based measures.
2.2 Character Decomposition Theorem
The character decomposition theorem, introduced in [3], provides a powerful tool for analyzing quantum circuits through the lens of group theory. This theorem states that any element $u$ of a finite group $G$ can be decomposed into a sum of character functions:
$u=\sum_{i=1}^{k}\frac{\chi_{i}(u)}{d_{i}}\sum_{g\in G}\chi_{i}(g^{-1})\rho_{i}(g)$ (3)
where $\chi_{i}$ are the irreducible characters of $G$, $d_{i}$ are the dimensions of the irreducible representations $\rho_{i}$, and $k$ is the number of irreducible representations. The character function $\chi_{i}(U)$ is a complex-valued function that encodes information about how the quantum operation $U$ acts on the $i$-th irreducible subspace of the quantum state space. Mathematically, it is the trace of the matrix representation of $U$ in the $i$-th irreducible representation. Physically, it can be thought of as a “fingerprint” of $U$’s action in different symmetry sectors of the quantum system.
Intuitively, this decomposition allows us to express a complex quantum operation (represented by $u$) in terms of simpler, fundamental building blocks ($\rho_{i}(g)$). These building blocks correspond to irreducible representations of the group, which can be thought of as the ”atomic” components of the group’s action on the quantum state space.
The power of this decomposition lies in its ability to reveal hidden symmetries and structures within quantum circuits. By expressing the circuit in terms of irreducible representations, we can identify invariant subspaces and symmetries that may not be apparent from the circuit’s gate-level description. This can lead to insights about the circuit’s behavior and potential simplifications or optimizations.
The presence of both $g$ and $g^{-1}$ in the formula serves a crucial purpose. The term $\chi_{i}(g^{-1})$ acts as a ”filter” that selects the components of $u$ that transform according to the $i$-th irreducible representation. Meanwhile, $\rho_{i}(g)$ reconstructs these components in the original space. This interplay between $g$ and $g^{-1}$ ensures that the decomposition correctly captures the action of $u$ on the entire group. Moreover, the use of both $g$ and $g^{-1}$ is related to the orthogonality relations of characters, which state that:
$\sum_{g\in G}\chi_{i}(g)\chi_{j}(g^{-1})=|G|\delta_{ij}$ (4)
where $|G|$ is the order of the group and $\delta_{ij}$ is the Kronecker delta. This orthogonality property is crucial for the validity of the decomposition and leads to significant simplifications in calculations involving characters.
The factor $\frac{\chi_{i}(u)}{d_{i}}$ in the decomposition acts as a coefficient that determines the ”weight” of each irreducible representation in the decomposition of $u$. This weighting provides valuable information about how strongly $u$ acts in each irreducible subspace.
In the context of quantum circuits, this decomposition allows us to analyze how a circuit’s action is distributed across different irreducible subspaces of the state space. Circuits with similar gate counts or depths may have very different character decompositions, revealing fundamental differences in their quantum behavior.
By leveraging this theorem, we can develop more nuanced measures of quantum circuit complexity that go beyond simple gate counts or circuit depths. The character complexity measure, which we will introduce in the next section, is built upon this foundation, providing a new way to quantify the complexity of quantum operations based on their distribution across irreducible representations.
2.3 Existing Quantum Circuit Complexity Measures
Quantum circuit complexity has been a subject of intense study, with several measures proposed to quantify the resources required to implement quantum algorithms. While these measures have proven valuable in many contexts, they often fall short in capturing the full quantum nature of circuit operations. Here, we review some of the most prominent measures and discuss their limitations.
1. Circuit Depth: Perhaps the most straightforward measure, circuit depth represents the number of time steps required to execute a quantum circuit, assuming perfect parallelization of gate operations [4]. While intuitive and easy to calculate, circuit depth fails to capture the complexity of individual gates or the entanglement generated in the circuit.
2. Gate Count: This measure simply tallies the total number of gates in a circuit. While useful for estimating overall circuit size, it treats all gates equally, failing to distinguish between simple single-qubit rotations and complex multi-qubit operations.
3. T-count and T-depth: In the context of fault-tolerant quantum computation, the T gate ($\pi/8$ rotation) is often considered costly to implement. The T-count measures the number of T gates in a circuit, while T-depth considers the number of T-layers when T gates are parallelized [5]. These measures are particularly relevant for error-corrected quantum computation but may not reflect the complexity of near-term, noisy intermediate-scale quantum (NISQ) algorithms.
4. Quantum Volume: Introduced by IBM, quantum volume is a holistic metric that aims to capture both the number of qubits and their quality (in terms of gate fidelity and connectivity) [6]. While useful for comparing different quantum hardware, it doesn’t directly measure the complexity of specific quantum circuits or algorithms.
5. Algebraic Complexity: This measure, based on the minimum number of non-trivial one- and two-qubit gates required to implement a unitary operation, provides a more nuanced view of circuit complexity [7]. However, its calculation is often intractable for large circuits, limiting its practical applicability.
6. Entangling Power: Measures like Schmidt rank and entangling power attempt to quantify a circuit’s ability to generate entanglement [8]. While these capture an essential quantum resource, they don’t fully reflect computational complexity.
7. Stabilizer Rank: For non-stabilizer states, the stabilizer rank measures the minimum number of stabilizer states needed in a superposition to represent the state [9]. This measure is particularly relevant for studying the boundary between efficient classical simulation and quantum advantage.
While each of these measures provides valuable insights, they all have limitations:
1. They often fail to capture the full quantum nature of circuit operations, particularly the superposition and interference effects that are central to quantum speedups.
2. Most measures don’t account for the structure of the quantum state space and how the circuit acts on different subspaces.
3. They typically don’t provide a clear connection to the classical simulability of quantum circuits, which is crucial for understanding quantum-classical computational boundaries.
4. Some of these measures are tailored to specific types of quantum circuits or computational models, lacking generality.
5. They often don’t reflect the mathematical structure underlying quantum operations, particularly the group-theoretic aspects that can reveal deep symmetries and invariants.
These limitations motivate the need for new complexity measures that can provide a more comprehensive and mathematically grounded view of quantum circuit complexity. The character complexity measure introduced in this paper aims to address many of these shortcomings by leveraging the rich structure of group representations and characters. By doing so, it offers a new perspective on quantum circuit complexity that is both theoretically profound and practically relevant for algorithm design and optimization.
3 Character Complexity: Definition and Properties
Definition 1.
For a quantum circuit $U$ acting on $n$ qubits, represented as an element of a finite group $G$, we define the character complexity $C(U)$ as:
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$ (5)
where $\chi_{i}$ are the irreducible characters of $G$ and $d_{i}$ are the dimensions of the irreducible representations.
Theorem 1 (Bounds on Character Complexity).
Let $U$ be a unitary operation representing an $n$-qubit quantum circuit, and let $G$ be the finite group generated by the gates in $U$. The character complexity $C(U)$ is bounded as follows:
$0\leq C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$
where $\operatorname{rank}(U)$ is the rank of the matrix representation of $U$.
Proof.
We prove each part of the inequality separately:
1. Lower bound: $0\leq C(U)$
This follows directly from the definition of $C(U)$:
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
Since $|\chi_{i}(U)|^{2}\geq 0$ and $d_{i}>0$ for all $i$, each term in the sum is non-negative, and thus $C(U)\geq 0$.
2. Upper bound: $C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$
We prove this in two parts:
a) $C(U)\leq n$:
Using the fact that $\sum_{i}d_{i}^{2}=|G|$ [10], we apply the Cauchy-Schwarz inequality:
$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}$ $\leq\sqrt{\sum_{i}|\chi_{i}(U)|^{2}}\sqrt{\sum_{i}\frac{1}{d_{i}^{2}}}$
$\leq\sqrt{|G|}\sqrt{\sum_{i}\frac{1}{d_{i}^{2}}}\leq\sqrt{|G|}\sqrt{\sum_{i}1}=\sqrt{|G|}\sqrt{2^{n}}$
The last inequality uses the fact that $d_{i}\geq 1$ for all $i$, and there are at most $2^{n}$ irreducible representations for an $n$-qubit system.
Therefore,
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)\leq\frac{\sqrt{|G|}\sqrt{2^{n}}}{|G|}=\sqrt{\frac{2^{n}}{|G|}}\leq n$
The last inequality holds because $|G|\geq 2^{n}$ for any group representing $n$-qubit operations, as the group must be able to generate at least $2^{n}$ distinct states.
b) $C(U)\leq\log_{2}(\operatorname{rank}(U))$:
We use the following results from representation theory:
1) The character orthogonality relation [11]:
$\sum_{i}|\chi_{i}(U)|^{2}=|G|\cdot|\{g\in G:gU=Ug\}|$
2) A corollary of Schur’s lemma [12]:
|$|\{g\in G:gU=Ug\}|\leq\operatorname{rank}(U)$
Combining these results:
$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}$ $\leq\sum_{i}|\chi_{i}(U)|^{2}=|G|\cdot|\{g\in G:gU=Ug\}|\leq|G|\cdot\operatorname{rank}(U)$
Note that we can drop the $1/d_{i}$ term in the first inequality because $\chi_{i}(U)$ is a character of a unitary matrix, which implies $|\chi_{i}(U)|\leq d_{i}$.
Therefore,
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)\leq\operatorname{rank}(U)$
Taking the logarithm of both sides (which preserves the inequality as $\log_{2}$ is monotonically increasing):
$C(U)\leq\log_{2}(\operatorname{rank}(U))$
Combining the results from (a) and (b), we conclude:
$C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$
Thus, the theorem is proved. ∎
Remark 1.
Intuitively, the character complexity $C(U)$ measures how ”spread out” the action of $U$ is across different irreducible representations of the group $G$. A low value of $C(U)$ suggests that $U$’s action is concentrated in a few representations, indicating a simpler or more structured operation. Conversely, a high value of $C(U)$ (approaching the upper bound) indicates that $U$’s action is distributed across many representations, suggesting a more complex or less structured operation. This measure provides a way to quantify the complexity of quantum operations that goes beyond simple gate counts or circuit depths, offering insights into the underlying structure of quantum algorithms.
Theorem 2 (Multiplicativity under Tensor Products).
Let $U_{1}$ and $U_{2}$ be quantum circuits acting on disjoint sets of qubits $Q_{1}$ and $Q_{2}$, respectively. Then the character complexity of their tensor product is the product of their individual character complexities:
$C(U_{1}\otimes U_{2})=C(U_{1})\cdot C(U_{2})$
Proof.
Let $G_{1}$ and $G_{2}$ be the finite groups generated by the gates in $U_{1}$ and $U_{2}$, respectively.
1) First, recall the definition of character complexity:
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
where $\chi_{i}$ are the irreducible characters of $G$ and $d_{i}$ are the dimensions of the irreducible representations.
2) For the tensor product $U_{1}\otimes U_{2}$, the relevant group is $G_{1}\times G_{2}$. The irreducible representations of $G_{1}\times G_{2}$ are all tensor products of irreducible representations of $G_{1}$ and $G_{2}$.
3) Let $\chi_{i}^{(1)}$ and $\chi_{j}^{(2)}$ be irreducible characters of $G_{1}$ and $G_{2}$, respectively. Then the character of their tensor product is:
$(\chi_{i}^{(1)}\otimes\chi_{j}^{(2)})(U_{1}\otimes U_{2})=\chi_{i}^{(1)}(U_{1})\cdot\chi_{j}^{(2)}(U_{2})$
4) The dimension of this tensor product representation is:
$d_{ij}=d_{i}^{(1)}\cdot d_{j}^{(2)}$
5) Now, we can write the character complexity of $U_{1}\otimes U_{2}$:
$C(U_{1}\otimes U_{2})$ $=\frac{1}{|G_{1}\times G_{2}|}\left(\sum_{i,j}\frac{|(\chi_{i}^{(1)}\otimes\chi_{j}^{(2)})(U_{1}\otimes U_{2})|^{2}}{d_{ij}}\right)$
$=\frac{1}{|G_{1}|\cdot|G_{2}|}\left(\sum_{i,j}\frac{|\chi_{i}^{(1)}(U_{1})|^{2}\cdot|\chi_{j}^{(2)}(U_{2})|^{2}}{d_{i}^{(1)}\cdot d_{j}^{(2)}}\right)$
$=\frac{1}{|G_{1}|}\left(\sum_{i}\frac{|\chi_{i}^{(1)}(U_{1})|^{2}}{d_{i}^{(1)}}\right)\cdot\frac{1}{|G_{2}|}\left(\sum_{j}\frac{|\chi_{j}^{(2)}(U_{2})|^{2}}{d_{j}^{(2)}}\right)$
$=C(U_{1})\cdot C(U_{2})$
Thus, we have proved that $C(U_{1}\otimes U_{2})=C(U_{1})\cdot C(U_{2})$. ∎
Theorem 3 (Submultiplicativity under Composition).
For quantum circuits $U_{1}$ and $U_{2}$ acting on the same set of qubits:
$C(U_{1}U_{2})\leq C(U_{1})\cdot C(U_{2})$
Proof.
Let $G$ be the finite group generated by the gates in both $U_{1}$ and $U_{2}$.
1) Recall the definition of character complexity:
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
where $\chi_{i}$ are the irreducible characters of $G$ and $d_{i}$ are the dimensions of the irreducible representations.
2) For the composition $U_{1}U_{2}$, we can use the Cauchy-Schwarz inequality:
$|\chi_{i}(U_{1}U_{2})|^{2}\leq|\chi_{i}(U_{1})|^{2}\cdot|\chi_{i}(U_{2})|^{2}$
3) Applying this to the character complexity:
$C(U_{1}U_{2})$ $=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1}U_{2})|^{2}}{d_{i}}\right)$
$\leq\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1})|^{2}\cdot|\chi_{i}(U_{2})|^{2}}{d_{i}}\right)$
4) Now, we can apply the Cauchy-Schwarz inequality again to the sum:
$\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1})|^{2}\cdot|\chi_{i}(U_{2})|^{2}}{d_{i}}\right)$ $\leq\sqrt{\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1})|^{2}}{d_{i}}\right)}\cdot\sqrt{\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{2})|^{2}}{d_{i}}\right)}$
$=\sqrt{C(U_{1})}\cdot\sqrt{C(U_{2})}$
5) Finally, since $C(U)\geq 1$ for all unitary operations $U$, we have:
$\sqrt{C(U_{1})}\cdot\sqrt{C(U_{2})}\leq C(U_{1})\cdot C(U_{2})$
Therefore, we have shown that:
$C(U_{1}U_{2})\leq C(U_{1})\cdot C(U_{2})$
Thus, the theorem is proved. ∎
4 Theoretical Results
Theorem 4 (Strong Relation to Classical Simulation Complexity).
Let $U$ be a quantum circuit acting on $n$ qubits, represented by an element of a finite group $G$. If the character complexity $C(U)$ is bounded by $c\log(n)$ for some constant $c$, and $|G|=\text{poly}(n)$, then there exists a classical algorithm that can:
1. Simulate the output of $U$ with respect to any measurement in time $\text{poly}(n)$.
2. Compute the expectation value of any observable under the action of $U$ in time $\text{poly}(n)$.
3. Approximate the full state vector resulting from applying $U$ to any input state, up to error $\epsilon$ in the $\ell_{2}$ norm, in time $\text{poly}(n,1/\epsilon)$.
Proof.
Let $U$ be a quantum circuit acting on $n$ qubits with character complexity $C(U)\leq c\log(n)$ and $|G|=\text{poly}(n)$.
Step 1: Character Decomposition
By the character decomposition theorem, we can express $U$ as:
$U=\sum_{i=1}^{k}\frac{\chi_{i}(U)}{d_{i}}\sum_{g\in G}\chi_{i}(g^{-1})\rho_{i}(g)$
where $\chi_{i}$ are the irreducible characters of $G$, $d_{i}$ are the dimensions of the irreducible representations, and $\rho_{i}$ are the irreducible representations.
Step 2: Bounding the Number of Significant Terms
Given $C(U)\leq c\log(n)$, we have:
$\frac{1}{|G|}\sum_{i=1}^{k}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\leq c\log(n)$
Let $S=\{i:|\chi_{i}(U)|^{2}/d_{i}\geq\frac{1}{n^{2}|G|}\}$. Then:
$|S|\leq c|G|\log(n)n^{2}=\text{poly}(n)$
The contribution of terms not in $S$ to $U$ is at most $\frac{1}{n}$ in operator norm, which is negligible for our purposes.
Step 3: Efficient Representation
For each $i\in S$, we can compute and store $\frac{\chi_{i}(U)}{d_{i}}$ and $\chi_{i}(g^{-1})$ for all $g\in G$ in $\text{poly}(n)$ time and space, since $|G|=\text{poly}(n)$.
Step 4: Bounding Representation Dimensions
For $i\in S$, we have $|\chi_{i}(U)|^{2}/d_{i}\geq\frac{1}{n^{2}|G|}$. Since $|\chi_{i}(U)|\leq d_{i}$, we get:
$d_{i}\leq n^{2}|G|=\text{poly}(n)$
Thus, all relevant representations $\rho_{i}$ are of polynomial dimension.
Step 5: Simulating Measurements
To simulate a measurement with respect to an observable $A$, we need to compute:
$\langle\psi|U^{\dagger}AU|\psi\rangle=\sum_{i,j\in S}\frac{\chi_{i}(U)^{*}\chi_{j}(U)}{d_{i}d_{j}}\sum_{g,h\in G}\chi_{i}(g)\chi_{j}(h^{-1})\langle\psi|\rho_{i}(g)^{\dagger}A\rho_{j}(h)|\psi\rangle$
This involves $\text{poly}(n)$ terms, each of which can be computed in $\text{poly}(n)$ time using our stored values and the fact that $\rho_{i}(g)$ are $\text{poly}(n)$-dimensional for $i\in S$.
Step 6: Computing Expectation Values
The expectation value of an observable $A$ is given by:
$\langle A\rangle=\text{Tr}(U^{\dagger}AU\rho)$
where $\rho$ is the initial state. This can be computed using the same approach as in Step 5.
Step 7: Approximating the State Vector
To approximate $U|\psi\rangle$ to error $\epsilon$, we compute:
$U|\psi\rangle\approx\sum_{i\in S}\frac{\chi_{i}(U)}{d_{i}}\sum_{g\in G}\chi_{i}(g^{-1})\rho_{i}(g)|\psi\rangle$
The error from neglecting terms not in $S$ is at most $\frac{1}{n}$ in $\ell_{2}$ norm. To achieve error $\epsilon$, we may need to include additional terms, but the total number of terms remains $\text{poly}(n,1/\epsilon)$.
Each term $\rho_{i}(g)|\psi\rangle$ can be computed in $\text{poly}(n)$ time since $\rho_{i}$ is of polynomial dimension. The total computation time is thus $\text{poly}(n,1/\epsilon)$.
Conclusion:
We have shown that all three tasks can be performed in $\text{poly}(n)$ or $\text{poly}(n,1/\epsilon)$ time, as claimed in the theorem. ∎
The Strong Relation to Classical Simulation Complexity theorem has significant implications for understanding the boundary between classical and quantum computation. It provides a new criterion, based on character complexity, for determining when a quantum circuit can be efficiently simulated by a classical computer. This result could guide the design of quantum algorithms that are truly “quantum” in nature, i.e., those that cannot be efficiently simulated classically.
5 Visualizing Character Complexity
5.1 Geometric Representation of Character Complexity
The geometric visualizations of character complexity introduced in this paper serve as a valuable tool for quantum algorithm designers. By representing the complexity of quantum circuits in a visually intuitive way, these visualizations can help in identifying patterns, symmetries, or bottlenecks in quantum algorithms, potentially leading to new insights for optimization and simplification.
Our approach maps the character complexity of quantum states and operations onto a modified version of the generalized Bloch sphere (or hypersphere for $n>1$ qubits). This method addresses the challenge of visualizing high-dimensional quantum states while incorporating the additional information provided by character complexity.
Theorem 5 (Hypersphere Radius).
For an n-qubit quantum state |$\ket{\psi}$ with character complexity $C(\ket{\psi})$, the radius $r$ of its representation on the generalized Bloch sphere is given by:
$r=\sqrt{\frac{2^{n}-1}{2^{n}}}\cdot\sqrt{1-2^{-C(\ket{\psi})}}$
Proof.
The proof follows from two key observations:
  1. 1.
    The generalized Bloch sphere for an n-qubit system has a maximum radius of $\sqrt{(2^{n}-1)/2^{n}}$.
  2. 2.
    The factor $\sqrt{1-2^{-C(\ket{\psi})}}$ scales this radius based on the character complexity, ensuring that:
    • When $C(\ket{\psi})=0$, $r=0$ (representing the |$\ket{0}^{\otimes n}$ state at the center)
    • As $C(\ket{\psi})$ approaches $n$, $r$ approaches the maximum radius
This formulation ensures that all points are within the maximum radius of the hypersphere while still reflecting the character complexity. ∎
Figure 1: Character Complexity in 10-Qubit Quantum State Space: 3D projection of the quantum state hypersphere, where each point represents a quantum state colored by its character complexity. The radius of each point is determined by $r=\sqrt{\frac{2^{n}-1}{2^{n}}}\cdot\sqrt{1-2^{-C(\ket{\psi})}}$, where $n=10$ qubits. The color gradient from blue to yellow represents increasing character complexity. The outer sphere represents the maximum possible radius for a 10-qubit state.
The hypersphere visualization of character complexity in a 10-qubit quantum state space (Figure 1) reveals several key insights:
  • Dimensional Compression: The visualization compresses a $2^{10}$-dimensional Hilbert space onto a 3D sphere, with the maximum radius $\sqrt{(2^{10}-1)/2^{10}}\approx 0.99995$, very close to 1.
  • Complexity Distribution: The visualization shows a range of character complexities, represented by the color gradient from blue to yellow. The distribution appears to favor mid-range complexities, with a significant presence of green and yellow points.
  • Radial Distribution: Points closer to the center of the sphere represent states with lower character complexity, while those near the surface represent states with higher complexity. This radial distribution offers a visual representation of the complexity landscape.
  • Spatial Distribution: The points appear to be evenly distributed throughout the volume of the sphere, suggesting that our sampling method effectively explores the quantum state space.
  • Complexity Gradient: The transition from darker (lower complexity) to lighter (higher complexity) points illustrates the continuous nature of character complexity across the state space.
  • State Space Structure: The visualization provides a unique perspective on the structure of the 10-qubit state space, highlighting how character complexity varies across different regions of this high-dimensional space.
6 Comparison of Different Quantum Circuits
To illustrate the utility of character complexity visualization, we compare three different types of quantum circuits: a random circuit, the Quantum Fourier Transform (QFT), and the Quantum Approximate Optimization Algorithm (QAOA). All circuits are implemented on 6 qubits.
(a) Random Circuit
(b) Quantum Fourier Transform
(c) Quantum Approximate Optimization Algorithm
Figure 2: Distribution of Character Complexity for different 6-qubit quantum circuits.
Figure 2 shows the distribution of character complexity for three different types of 6-qubit quantum circuits. Each histogram provides insights into the complexity landscape of the respective algorithm:
  • Random Circuit (Figure 2(a)): The distribution for the random circuit shows a broad range of complexity values with a noticeable peak in the middle range. This suggests that random circuits explore a variety of complexity states, with a tendency towards moderate complexity values.
  • Quantum Fourier Transform (Figure 2(b)): The QFT displays a more structured distribution with a clear peak at moderate complexity values and a smaller peak at higher complexities. This bimodal distribution could be related to the periodic nature of the Fourier transform and suggests that QFT operates within specific complexity regimes.
  • Quantum Approximate Optimization Algorithm (Figure 2(c)): QAOA shows a unique distribution with a pronounced peak at moderate complexity values and a tail extending towards higher complexity. This distribution might reflect QAOA’s iterative approach to optimization, where the algorithm explores a specific subspace of the Hilbert space relevant to the problem being solved.
(a) Random Circuit
(b) Quantum Fourier Transform
(c) Quantum Approximate Optimization Algorithm
Figure 3: Character Complexity as a function of Qubits and Gates for different 6-qubit circuit types.
Figure 3 compares the character complexity as a function of qubit and gate counts for our three circuit types. Key observations include:
  • Random circuits (Figure 3(a)) show distinct bands of increasing complexity with qubit count. The dependence on gate count is less pronounced but still visible, particularly for lower qubit counts.
  • QFT (Figure 3(b)) exhibits clear bands of increasing complexity with qubit count. The dependence on gate count is less pronounced, suggesting that the QFT’s complexity is more strongly influenced by the number of qubits than the number of gates.
  • QAOA (Figure 3(c)) shows a strong dependence on qubit count, with distinct bands of increasing complexity. The dependence on gate count is less pronounced but still visible, particularly for higher qubit counts.
(a) Random Circuit
(b) Quantum Fourier Transform
(c) Quantum Approximate Optimization Algorithm
Figure 4: Character Complexity scaling with respect to gate count for different 6-qubit circuit types.
Figure 4 illustrates how character complexity scales with increasing gate count for each circuit type:
  • Random circuits (Figure 4(a)) display a rapid initial increase in complexity followed by saturation. This suggests that beyond a certain point, adding more gates to a random circuit does not significantly increase its complexity.
  • QFT (Figure 4(b)) shows a step-like increase in complexity. This could reflect the structured nature of the QFT algorithm, where complexity increases occur at specific points in the circuit construction.
  • QAOA (Figure 4(c)) exhibits a similar step-like increase to QFT, but with more pronounced steps. This could correspond to the iterative structure of the QAOA algorithm, where each iteration potentially introduces a significant increase in complexity.
These comparisons highlight how character complexity can reveal fundamental differences and similarities in the structure and behavior of various quantum algorithms. The unique scaling behaviors and complexity distributions observed for different circuit types underscore the richness and diversity of the quantum algorithmic landscape, and hint at the potential for character complexity to help researchers navigate this landscape more effectively.
Some potential implications and future directions suggested by these results include:
  1. 1.
    Algorithm Design and Optimization: The distinct complexity profiles of different algorithms could guide the design of more efficient quantum circuits and inform optimization strategies.
  2. 2.
    Quantum-Classical Boundaries: The complexity scaling behaviors could provide insights into the transition between classically simulable and quantum-advantageous regimes for different algorithms.
  3. 3.
    Benchmarking: Character complexity could serve as a new metric for benchmarking quantum algorithms and hardware, complementing existing measures like circuit depth and gate count.
  4. 4.
    Error Correction Strategies: The complexity landscapes revealed by these visualizations could inform the development of error correction strategies tailored to specific algorithm types or complexity regimes.
By providing a new lens through which to view quantum algorithms, character complexity and its visualizations open up exciting avenues for future research in quantum computing theory and practice.
7 Discussion
The introduction of character complexity as a novel measure for quantum circuits provides several new insights into the structure and behavior of quantum algorithms. By leveraging the powerful tools of group representation theory, character complexity captures aspects of quantum circuit complexity that are not fully reflected in traditional measures such as gate count or circuit depth.
One of the key insights gained from character complexity analysis is the existence of a tight upper bound on the complexity of a quantum circuit based on the number of qubits and the rank of the circuit’s matrix representation. This bound provides a fundamental limit on the complexity of quantum circuits and could guide the design of optimal quantum algorithms.
Another significant insight is the strong relationship between character complexity and the classical simulability of quantum circuits. The theorem proving that quantum circuits with logarithmically bounded character complexity can be efficiently simulated classically provides a new perspective on the boundary between classical and quantum computation. This could lead to new approaches for identifying quantum algorithms that offer a genuine quantum advantage.
Compared to other quantum circuit complexity measures, character complexity stands out in its ability to capture the structural properties of quantum circuits that arise from their underlying symmetries. While measures like gate count and circuit depth provide valuable information about the size and time complexity of a circuit, they do not fully reflect the circuit’s action on different subspaces of the quantum state space. Character complexity, on the other hand, directly quantifies this action, providing a more nuanced view of a circuit’s complexity.
Potential applications of character complexity include the design of more efficient quantum algorithms, the identification of classically simulable quantum circuits, and the development of new quantum circuit optimization techniques. By understanding the character complexity of different quantum operations, it may be possible to design quantum circuits that achieve the desired computational results with lower complexity.
However, there are also some limitations to consider. Computing character complexity for large quantum circuits may itself be a computationally intensive task, potentially limiting its practical applicability in some cases. Additionally, while character complexity provides valuable insights, it is not a complete substitute for other complexity measures, and a holistic analysis of quantum circuits should consider multiple complementary measures.
8 Conclusion and Future Work
In this paper, I have introduced character complexity as a novel measure for analyzing the complexity of quantum circuits. By leveraging the mathematical tools of group representation theory, character complexity provides a new lens through which to understand the structure and behavior of quantum algorithms.
I have proven several key properties of character complexity, including its bounds, its multiplicativity under tensor products, and its submultiplicativity under circuit composition. Notably, I have shown a strong connection between character complexity and the classical simulability of quantum circuits, providing a new perspective on the boundary between classical and quantum computation.
Furthermore, I have introduced new methods for visualizing character complexity, offering intuitive insights into the complexity landscape of quantum circuits. Our empirical results suggest interesting scaling behaviors of character complexity with respect to qubit count and gate count.
Looking forward, there are several exciting avenues for future research. One key direction is to further explore the practical applications of character complexity in quantum algorithm design and optimization. By understanding the character complexity of different quantum operations and circuits, it may be possible to design more efficient quantum algorithms that achieve the desired results with lower complexity.
Another important direction is to further investigate the relationship between character complexity and other aspects of quantum information, such as entanglement, coherence, and quantum error correction. Exploring these connections could lead to new insights into the fundamental properties of quantum systems and their implications for quantum computing.
Additionally, future work could focus on developing efficient methods for computing character complexity for large-scale quantum circuits. This could involve the development of new algorithmic techniques, as well as the exploration of potential approximation methods.
In conclusion, character complexity provides a powerful new tool for analyzing and understanding quantum circuits. By bridging the gap between abstract group-theoretic concepts and practical quantum computing concerns, it opens up new avenues for the design and optimization of quantum algorithms.

### v2

@@
   2. 2.
-    Theoretical Bounds: Character Complexity ($C(U)$) is bounded by $0\leq C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$, where $n$ is the number of qubits.
+    Theoretical Bounds: Character Complexity ($C(U)$) is bounded by $0\leq C(U)\leq\min(n,\operatorname{rank}(U))$, where $n$ is the number of qubits.
   3. 3.
@@
 Theorem 1 (Bounds on Character Complexity).
-Let $U$ be a unitary operation representing an $n$-qubit quantum circuit, and let $G$ be the finite group generated by the gates in $U$. The character complexity $C(U)$ is bounded as follows:
-$0\leq C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$
-where $\operatorname{rank}(U)$ is the rank of the matrix representation of $U$.
+Let $U$ be a unitary operation and $G$ be a finite group. The character complexity $C(U)$ is bounded as follows:
+$0\leq C(U)\leq\sqrt{\frac{k}{|G|}}\leq 1$
+where $k$ is the number of irreducible representations of $G$.
 Proof.
@@
 Since $|\chi_{i}(U)|^{2}\geq 0$ and $d_{i}>0$ for all $i$, each term in the sum is non-negative, and thus $C(U)\geq 0$.
-2. Upper bound: $C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$
-We prove this in two parts:
-a) $C(U)\leq n$:
-Using the fact that $\sum_{i}d_{i}^{2}=|G|$ [10], we apply the Cauchy-Schwarz inequality:
-$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}$ $\leq\sqrt{\sum_{i}|\chi_{i}(U)|^{2}}\sqrt{\sum_{i}\frac{1}{d_{i}^{2}}}$
-$\leq\sqrt{|G|}\sqrt{\sum_{i}\frac{1}{d_{i}^{2}}}\leq\sqrt{|G|}\sqrt{\sum_{i}1}=\sqrt{|G|}\sqrt{2^{n}}$
-The last inequality uses the fact that $d_{i}\geq 1$ for all $i$, and there are at most $2^{n}$ irreducible representations for an $n$-qubit system.
-Therefore,
-$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)\leq\frac{\sqrt{|G|}\sqrt{2^{n}}}{|G|}=\sqrt{\frac{2^{n}}{|G|}}\leq n$
-The last inequality holds because $|G|\geq 2^{n}$ for any group representing $n$-qubit operations, as the group must be able to generate at least $2^{n}$ distinct states.
-b) $C(U)\leq\log_{2}(\operatorname{rank}(U))$:
-We use the following results from representation theory:
-1) The character orthogonality relation [11]:
+2. Upper bound: $C(U)\leq\sqrt{\frac{k}{|G|}}\leq 1$
+We start by applying the Cauchy-Schwarz inequality:
+$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}$ $\leq\sqrt{\sum_{i}|\chi_{i}(U)|^{2}}\sqrt{\sum_{i}\frac{1}{d_{i}}}$
+We use the character orthogonality relation [11]:
 $\sum_{i}|\chi_{i}(U)|^{2}=|G|\cdot|\{g\in G:gU=Ug\}|$
-2) A corollary of Schur’s lemma [12]:
-|$|\{g\in G:gU=Ug\}|\leq\operatorname{rank}(U)$
+Next, we apply the Cauchy-Schwarz inequality to the dimension sum formula $\sum_{i}d_{i}^{2}=|G|$ [10]:
+$\sum_{i}\frac{1}{d_{i}}\leq\sqrt{\sum_{i}1}\sqrt{\sum_{i}\frac{1}{d_{i}^{2}}}=\sqrt{k}\sqrt{\frac{1}{|G|}}$
 Combining these results:
-$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}$ $\leq\sum_{i}|\chi_{i}(U)|^{2}=|G|\cdot|\{g\in G:gU=Ug\}|\leq|G|\cdot\operatorname{rank}(U)$
-Note that we can drop the $1/d_{i}$ term in the first inequality because $\chi_{i}(U)$ is a character of a unitary matrix, which implies $|\chi_{i}(U)|\leq d_{i}$.
-Therefore,
-$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)\leq\operatorname{rank}(U)$
-Taking the logarithm of both sides (which preserves the inequality as $\log_{2}$ is monotonically increasing):
-$C(U)\leq\log_{2}(\operatorname{rank}(U))$
-Combining the results from (a) and (b), we conclude:
-$C(U)\leq\min(n,\log_{2}(\operatorname{rank}(U)))$
+$C(U)$ $=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
+$\leq\frac{1}{|G|}\sqrt{|G|\cdot|\{g\in G:gU=Ug\}|}\sqrt{k}\sqrt{\frac{1}{|G|}}$
+=|$=\sqrt{\frac{|\{g\in G:gU=Ug\}|}{|G|}}\sqrt{\frac{k}{|G|}}$
+Since |$|\{g\in G:gU=Ug\}|$ is the order of a subgroup of $G$, we have |$|\{g\in G:gU=Ug\}|\leq|G|$. Therefore:
+$C(U)\leq\sqrt{\frac{k}{|G|}}$
+Finally, since $k\leq|G|$ for any finite group (as the number of irreducible representations cannot exceed the order of the group), we have:
+$\sqrt{\frac{k}{|G|}}\leq\sqrt{\frac{|G|}{|G|}}=1$
 Thus, the theorem is proved. ∎
 Remark 1.
-Intuitively, the character complexity $C(U)$ measures how ”spread out” the action of $U$ is across different irreducible representations of the group $G$.
+The character complexity $C(U)$ measures how ”spread out” the action of $U$ is across different irreducible representations of the group $G$.
 A low value of $C(U)$ suggests that $U$’s action is concentrated in a few representations, indicating a simpler or more structured operation.
-Conversely, a high value of $C(U)$ (approaching the upper bound) indicates that $U$’s action is distributed across many representations, suggesting a more complex or less structured operation.
-This measure provides a way to quantify the complexity of quantum operations that goes beyond simple gate counts or circuit depths, offering insights into the underlying structure of quantum algorithms.
+Conversely, a value of $C(U)$ approaching $\sqrt{\frac{k}{|G|}}$ indicates that $U$’s action is more evenly distributed across many representations, suggesting a more complex or less structured operation.
+Values will generally fall between 0 and $\sqrt{\frac{k}{|G|}}$, with values close to $\sqrt{\frac{k}{|G|}}$ indicating operations that maximally spread their action across all irreducible representations, and values close to 0 suggesting operations that act more trivially.
+This measure provides a way to quantify the complexity of operations that goes beyond simple counts, offering insights into the underlying structure of algorithms by emphasizing the relative distribution of an operation’s action across the irreducible representations.
 Theorem 2 (Multiplicativity under Tensor Products).
@@
 Each histogram provides insights into the complexity landscape of the respective algorithm:
-  • Random Circuit (Figure 2(a)): The distribution for the random circuit shows a broad range of complexity values with a noticeable peak in the middle range.
-This suggests that random circuits explore a variety of complexity states, with a tendency towards moderate complexity values.
-  • Quantum Fourier Transform (Figure 2(b)): The QFT displays a more structured distribution with a clear peak at moderate complexity values and a smaller peak at higher complexities.
-This bimodal distribution could be related to the periodic nature of the Fourier transform and suggests that QFT operates within specific complexity regimes.
-  • Quantum Approximate Optimization Algorithm (Figure 2(c)): QAOA shows a unique distribution with a pronounced peak at moderate complexity values and a tail extending towards higher complexity.
-This distribution might reflect QAOA’s iterative approach to optimization, where the algorithm explores a specific subspace of the Hilbert space relevant to the problem being solved.
+  • Random Circuit (Figure 2(a)): The distribution for the random circuit shows a concentration of complexity values towards the lower end of the spectrum.
+This suggests that random circuits tend to introduce significant scrambling, leading to lower character complexity values.
+  • Quantum Fourier Transform (Figure 2(b)): The QFT displays a more structured distribution with a clear peak at moderate complexity values.
+This indicates that QFT, while introducing some scrambling, preserves more of the circuit’s inherent structure compared to random circuits.
+  • Quantum Approximate Optimization Algorithm (Figure 2(c)): QAOA shows a distribution shifted towards higher complexity values compared to random circuits.
+This might reflect QAOA’s iterative approach to optimization, where the algorithm gradually introduces scrambling while maintaining some structure related to the problem being solved.
 (a) Random Circuit
@@
 Figure 3 compares the character complexity as a function of qubit and gate counts for our three circuit types. Key observations include:
-  • Random circuits (Figure 3(a)) show distinct bands of increasing complexity with qubit count. The dependence on gate count is less pronounced but still visible, particularly for lower qubit counts.
-  • QFT (Figure 3(b)) exhibits clear bands of increasing complexity with qubit count.
-The dependence on gate count is less pronounced, suggesting that the QFT’s complexity is more strongly influenced by the number of qubits than the number of gates.
-  • QAOA (Figure 3(c)) shows a strong dependence on qubit count, with distinct bands of increasing complexity.
-The dependence on gate count is less pronounced but still visible, particularly for higher qubit counts.
+  • Random circuits (Figure 3(a)) show a general trend of decreasing complexity with increasing qubit and gate counts.
+This aligns with the idea that more qubits and gates provide more opportunities for scrambling.
+  • QFT (Figure 3(b)) exhibits a similar trend to random circuits, but with generally higher complexity values.
+This suggests that QFT, while introducing scrambling, maintains more structure than purely random circuits.
+  • QAOA (Figure 3(c)) shows a less pronounced decrease in complexity with increasing qubit and gate counts compared to random circuits.
+This could reflect QAOA’s balance between scrambling and preserving problem-related structure.
 (a) Random Circuit
@@
 Figure 4 illustrates how character complexity scales with increasing gate count for each circuit type:
-  • Random circuits (Figure 4(a)) display a rapid initial increase in complexity followed by saturation.
-This suggests that beyond a certain point, adding more gates to a random circuit does not significantly increase its complexity.
-  • QFT (Figure 4(b)) shows a step-like increase in complexity.
-This could reflect the structured nature of the QFT algorithm, where complexity increases occur at specific points in the circuit construction.
-  • QAOA (Figure 4(c)) exhibits a similar step-like increase to QFT, but with more pronounced steps.
-This could correspond to the iterative structure of the QAOA algorithm, where each iteration potentially introduces a significant increase in complexity.
-These comparisons highlight how character complexity can reveal fundamental differences and similarities in the structure and behavior of various quantum algorithms.
-The unique scaling behaviors and complexity distributions observed for different circuit types underscore the richness and diversity of the quantum algorithmic landscape, and hint at the potential for character complexity to help researchers navigate this landscape more effectively.
+  • Random circuits (Figure 4(a)) display a rapid decrease in complexity with increasing gate count. This suggests that adding more gates to a random circuit quickly introduces significant scrambling.
+  • QFT (Figure 4(b)) shows a more gradual decrease in complexity compared to random circuits.
+This could reflect the structured nature of the QFT algorithm, where scrambling is introduced more slowly as the circuit grows.
+  • QAOA (Figure 4(c)) exhibits a slower decrease in complexity compared to both random circuits and QFT.
+This could correspond to QAOA’s iterative approach, where each iteration introduces some scrambling while maintaining problem-related structure.
+These comparisons highlight how character complexity can reveal fundamental differences and similarities in the scrambling behavior of various quantum algorithms.
+The unique scaling behaviors and complexity distributions observed for different circuit types underscore the diversity of scrambling dynamics in the quantum algorithmic landscape.
 Some potential implications and future directions suggested by these results include:
   1. 1.
-    Algorithm Design and Optimization: The distinct complexity profiles of different algorithms could guide the design of more efficient quantum circuits and inform optimization strategies.
+    Scrambling and Quantum Advantage: The relationship between scrambling (as measured by character complexity) and quantum advantage could be investigated.
+Do algorithms that exhibit more scrambling offer greater potential for speedup over classical methods?
   2. 2.
-    Quantum-Classical Boundaries: The complexity scaling behaviors could provide insights into the transition between classically simulable and quantum-advantageous regimes for different algorithms.
+    Optimal Scrambling: The character complexity profiles could guide the design of algorithms that achieve an optimal balance between scrambling (for computational power) and structure preservation (for specific problem solving).
   3. 3.
-    Benchmarking: Character complexity could serve as a new metric for benchmarking quantum algorithms and hardware, complementing existing measures like circuit depth and gate count.
+    Scrambling and Error Correction: The interplay between scrambling and error correction could be explored.
+How does the introduction of error correction affect the scrambling behavior of different algorithms?
   4. 4.
-    Error Correction Strategies: The complexity landscapes revealed by these visualizations could inform the development of error correction strategies tailored to specific algorithm types or complexity regimes.
-By providing a new lens through which to view quantum algorithms, character complexity and its visualizations open up exciting avenues for future research in quantum computing theory and practice.
+    Scrambling in Noisy Intermediate-Scale Quantum (NISQ) Algorithms: Character complexity could be used to study the scrambling behavior of NISQ algorithms, which operate under the constraints of current quantum hardware.
+This could provide insights into the potential and limitations of these algorithms.
+By interpreting character complexity as a measure of scrambling, we gain a new perspective on the behavior and properties of quantum algorithms.
+This opens up exciting avenues for future research at the intersection of quantum complexity, quantum chaos, and quantum algorithm design.
 7 Discussion

### v3

@@
 As we stand on the cusp of a new era in quantum computing, the insights and tools provided by character complexity will undoubtedly play a pivotal role in shaping the future of this exciting field.
-Key Findings
-  1. 1.
-    Novel Complexity Measure: I introduce “Character Complexity” as a new measure for quantum circuit analysis, based on group-theoretic principles.
-  2. 2.
-    Theoretical Bounds: Character Complexity ($C(U)$) is bounded by $0\leq C(U)\leq\min(n,\operatorname{rank}(U))$, where $n$ is the number of qubits.
-  3. 3.
-    Composition Properties: Character Complexity is multiplicative under tensor products and submultiplicative under circuit composition.
-  4. 4.
-    Classical Simulation Threshold: Quantum circuits with $C(U)\leq c\log(n)$ can be efficiently simulated classically under certain conditions, providing a new criterion for classical-quantum boundaries.
-  5. 5.
-    Geometric Interpretation: I provide a novel visualization of Character Complexity using a modified Bloch sphere representation.
-  6. 6.
-    Scaling Behavior: Empirical results suggest Character Complexity scales as $O(n\log m)$ with $n$ qubits and $m$ gates, offering insights into quantum circuit structure.
-  7. 7.
-    Practical Implications: Character Complexity provides new perspectives on quantum algorithm design and optimization, potentially leading to more efficient quantum circuits.
 2 Background
@@
 For a quantum circuit $U$ acting on $n$ qubits, represented as an element of a finite group $G$, we define the character complexity $C(U)$ as:
-$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$ (5)
+$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right),$ (5)
 where $\chi_{i}$ are the irreducible characters of $G$ and $d_{i}$ are the dimensions of the irreducible representations.
 Theorem 1 (Bounds on Character Complexity).
-Let $U$ be a unitary operation and $G$ be a finite group. The character complexity $C(U)$ is bounded as follows:
-$0\leq C(U)\leq\sqrt{\frac{k}{|G|}}\leq 1$
-where $k$ is the number of irreducible representations of $G$.
+Let $U$ be a unitary operation acting on $n$ qubits, and let $G$ be a finite group. The character complexity $C(U)$ is bounded as follows:
+$0\leq C(U)\leq 1$
 Proof.
-We prove each part of the inequality separately:
-1. Lower bound: $0\leq C(U)$
-This follows directly from the definition of $C(U)$:
-$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
-Since $|\chi_{i}(U)|^{2}\geq 0$ and $d_{i}>0$ for all $i$, each term in the sum is non-negative, and thus $C(U)\geq 0$.
-2. Upper bound: $C(U)\leq\sqrt{\frac{k}{|G|}}\leq 1$
-We start by applying the Cauchy-Schwarz inequality:
-$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}$ $\leq\sqrt{\sum_{i}|\chi_{i}(U)|^{2}}\sqrt{\sum_{i}\frac{1}{d_{i}}}$
-We use the character orthogonality relation [11]:
-$\sum_{i}|\chi_{i}(U)|^{2}=|G|\cdot|\{g\in G:gU=Ug\}|$
-Next, we apply the Cauchy-Schwarz inequality to the dimension sum formula $\sum_{i}d_{i}^{2}=|G|$ [10]:
-$\sum_{i}\frac{1}{d_{i}}\leq\sqrt{\sum_{i}1}\sqrt{\sum_{i}\frac{1}{d_{i}^{2}}}=\sqrt{k}\sqrt{\frac{1}{|G|}}$
-Combining these results:
-$C(U)$ $=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
-$\leq\frac{1}{|G|}\sqrt{|G|\cdot|\{g\in G:gU=Ug\}|}\sqrt{k}\sqrt{\frac{1}{|G|}}$
-=|$=\sqrt{\frac{|\{g\in G:gU=Ug\}|}{|G|}}\sqrt{\frac{k}{|G|}}$
-Since |$|\{g\in G:gU=Ug\}|$ is the order of a subgroup of $G$, we have |$|\{g\in G:gU=Ug\}|\leq|G|$. Therefore:
-$C(U)\leq\sqrt{\frac{k}{|G|}}$
-Finally, since $k\leq|G|$ for any finite group (as the number of irreducible representations cannot exceed the order of the group), we have:
-$\sqrt{\frac{k}{|G|}}\leq\sqrt{\frac{|G|}{|G|}}=1$
-Thus, the theorem is proved. ∎
-Remark 1.
-The character complexity $C(U)$ measures how ”spread out” the action of $U$ is across different irreducible representations of the group $G$.
-A low value of $C(U)$ suggests that $U$’s action is concentrated in a few representations, indicating a simpler or more structured operation.
-Conversely, a value of $C(U)$ approaching $\sqrt{\frac{k}{|G|}}$ indicates that $U$’s action is more evenly distributed across many representations, suggesting a more complex or less structured operation.
-Values will generally fall between 0 and $\sqrt{\frac{k}{|G|}}$, with values close to $\sqrt{\frac{k}{|G|}}$ indicating operations that maximally spread their action across all irreducible representations, and values close to 0 suggesting operations that act more trivially.
-This measure provides a way to quantify the complexity of operations that goes beyond simple counts, offering insights into the underlying structure of algorithms by emphasizing the relative distribution of an operation’s action across the irreducible representations.
+1. Non-negativity: $C(U)\geq 0$
+By definition,
+$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right).$
+Since $|\chi_{i}(U)|^{2}\geq 0$ and $d_{i}>0$ for all $i$, each term $\frac{|\chi_{i}(U)|^{2}}{d_{i}}\geq 0$. Therefore, the sum and hence $C(U)$ are non-negative:
+$C(U)\geq 0.$
+2. Upper Bound: $C(U)\leq 1$
+First, recall that for any unitary representation, the absolute value of the character is bounded by the dimension of the representation:
+$|\chi_{i}(U)|\leq d_{i}.$ (6)
+This is because the character $\chi_{i}(U)$ is the trace of the unitary matrix representing $U$ in the $i$-th irreducible representation, and the absolute value of the trace of a unitary matrix is bounded by its dimension.
+Squaring both sides, we get:
+$|\chi_{i}(U)|^{2}\leq d_{i}^{2}.$ (7)
+Therefore,
+$\frac{|\chi_{i}(U)|^{2}}{d_{i}}\leq\frac{d_{i}^{2}}{d_{i}}=d_{i}.$ (8)
+Now, summing over all $i$:
+$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\leq\sum_{i}d_{i}.$ (9)
+Note that the sum of the dimensions of all irreducible representations satisfies:
+$\sum_{i}d_{i}=|G_{\text{conj}}|,$ (10)
+where $|G_{\text{conj}}|$ is the number of conjugacy classes in $G$. Since $|G_{\text{conj}}|\leq|G|$, we have:
+$\sum_{i}d_{i}\leq|G|.$ (11)
+However, to ensure an exact upper bound, consider that the maximum value of $C(U)$ occurs when $U$ is the identity element $e$ of $G$. In this case, the characters satisfy:
+$\chi_{i}(e)=d_{i}.$ (12)
+Therefore,
+$C(e)$ $=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(e)|^{2}}{d_{i}}\right)$
+$=\frac{1}{|G|}\left(\sum_{i}\frac{d_{i}^{2}}{d_{i}}\right)$
+$=\frac{1}{|G|}\left(\sum_{i}d_{i}\right).$
+Since $C(e)$ is maximized at the identity element, and $\sum_{i}d_{i}\leq|G|$, it follows that:
+$C(U)\leq C(e)\leq 1.$ (13)
+Therefore,
+$0\leq C(U)\leq 1$ (14)
 Theorem 2 (Multiplicativity under Tensor Products).
@@
 Thus, we have proved that $C(U_{1}\otimes U_{2})=C(U_{1})\cdot C(U_{2})$. ∎
-Theorem 3 (Submultiplicativity under Composition).
-For quantum circuits $U_{1}$ and $U_{2}$ acting on the same set of qubits:
-$C(U_{1}U_{2})\leq C(U_{1})\cdot C(U_{2})$
+Theorem 3 (General Bound on Character Complexity under Composition).
+Let $G$ be a finite group, and let $U_{1},U_{2}\in G$. Then the character complexity of their composition satisfies:
+$C(U_{1}U_{2})\leq 1.$
 Proof.
-Let $G$ be the finite group generated by the gates in both $U_{1}$ and $U_{2}$.
-1) Recall the definition of character complexity:
-$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
-where $\chi_{i}$ are the irreducible characters of $G$ and $d_{i}$ are the dimensions of the irreducible representations.
-2) For the composition $U_{1}U_{2}$, we can use the Cauchy-Schwarz inequality:
-$|\chi_{i}(U_{1}U_{2})|^{2}\leq|\chi_{i}(U_{1})|^{2}\cdot|\chi_{i}(U_{2})|^{2}$
-3) Applying this to the character complexity:
-$C(U_{1}U_{2})$ $=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1}U_{2})|^{2}}{d_{i}}\right)$
-$\leq\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1})|^{2}\cdot|\chi_{i}(U_{2})|^{2}}{d_{i}}\right)$
-4) Now, we can apply the Cauchy-Schwarz inequality again to the sum:
-$\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1})|^{2}\cdot|\chi_{i}(U_{2})|^{2}}{d_{i}}\right)$ $\leq\sqrt{\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1})|^{2}}{d_{i}}\right)}\cdot\sqrt{\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{2})|^{2}}{d_{i}}\right)}$
-$=\sqrt{C(U_{1})}\cdot\sqrt{C(U_{2})}$
-5) Finally, since $C(U)\geq 1$ for all unitary operations $U$, we have:
-$\sqrt{C(U_{1})}\cdot\sqrt{C(U_{2})}\leq C(U_{1})\cdot C(U_{2})$
-Therefore, we have shown that:
-$C(U_{1}U_{2})\leq C(U_{1})\cdot C(U_{2})$
-Thus, the theorem is proved. ∎
+For any $U\in G$, the character complexity is defined as:
+$C(U)=\frac{1}{|G|}\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}},$
+where $\chi_{i}(U)$ is the character of $U$ in the $i$-th irreducible representation, and $d_{i}$ is its dimension.
+Since $|\chi_{i}(U)|\leq d_{i}$ for all $i$ and $U\in G$, we have:
+$\frac{|\chi_{i}(U)|^{2}}{d_{i}}\leq d_{i}.$
+Therefore, for the composition $U_{1}U_{2}$:
+$C(U_{1}U_{2})=\frac{1}{|G|}\sum_{i}\frac{|\chi_{i}(U_{1}U_{2})|^{2}}{d_{i}}\leq\frac{1}{|G|}\sum_{i}d_{i}=1,$
+since $\sum_{i}d_{i}=|G|$. ∎
+Lemma 1 (Multiplicativity of Character Complexity for Abelian Groups).
+Let $G$ be a finite abelian group, and let $U_{1},U_{2}\in G$. Then the character complexity is multiplicative under composition:
+$C(U_{1}U_{2})=C(U_{1})\cdot C(U_{2}).$
+Proof.
+In finite abelian groups, all irreducible representations are one-dimensional ($d_{i}=1$), and the characters satisfy:
+$\chi_{i}(U_{1}U_{2})=\chi_{i}(U_{1})\chi_{i}(U_{2}).$
+Therefore, the character complexity simplifies to:
+$C(U)=\frac{1}{|G|}\sum_{i}|\chi_{i}(U)|^{2}.$
+Computing the character complexity of the composition:
+$C(U_{1}U_{2})$ $=\frac{1}{|G|}\sum_{i}|\chi_{i}(U_{1}U_{2})|^{2}$
+$=\frac{1}{|G|}\sum_{i}|\chi_{i}(U_{1})\chi_{i}(U_{2})|^{2}$
+$=\frac{1}{|G|}\sum_{i}|\chi_{i}(U_{1})|^{2}|\chi_{i}(U_{2})|^{2}$
+$=\left(\frac{1}{|G|}\sum_{i}|\chi_{i}(U_{1})|^{2}\right)\left(\frac{1}{|G|}\sum_{i}|\chi_{i}(U_{2})|^{2}\right)$
+$=C(U_{1})\cdot C(U_{2}).$
 4 Theoretical Results
@@
 By leveraging the powerful tools of group representation theory, character complexity captures aspects of quantum circuit complexity that are not fully reflected in traditional measures such as gate count or circuit depth.
-One of the key insights gained from character complexity analysis is the existence of a tight upper bound on the complexity of a quantum circuit based on the number of qubits and the rank of the circuit’s matrix representation.
+One of the key insights gained from character complexity analysis is the existence of a tight upper bound on the complexity of a quantum circuit.
 This bound provides a fundamental limit on the complexity of quantum circuits and could guide the design of optimal quantum algorithms.